OpenStudy (anonymous):

Zoe keeps track of the miles per gallon her car gets per week. She has accumulated the following data: (1, 24), (2, 24.38), (3, 24.76), (4, 25.14) What is the common difference or ratio? The common difference is 0.38. The common difference is 0.76. The common ratio is 0.99. The common ratio is 1.02. just need the answer

OpenStudy (anonymous):

yo help

OpenStudy (anonymous):

someone

OpenStudy (jdoe0001):

\({\bf (1, 24), (2, 24.38), (3, 24.76), (4, 25.14)}\implies \begin{array}{ccccll} 1&2&3&4\\ \hline\\ 24&24.38&24.76&25.14 \end{array}\) what do you think?

OpenStudy (anonymous):

C

OpenStudy (jdoe0001):

hmmm why do you think is "c" anyway?

OpenStudy (anonymous):

educational guess

OpenStudy (jdoe0001):

heheeh, you mean "wild guess" so... do you know what's meant by \(\bf \text{common difference or ratio?}\)

OpenStudy (anonymous):

no jaja

OpenStudy (jdoe0001):

common difference, the amount the next term differs from the previous one say 2, 4 ,6 <-- common difference? well, 4 is previous + 2, so 2+2 =4 6 is previous + 2, so 4+2 = 6 so, the next term is adding 2 from the previous thus the common difference would be "2" on that

OpenStudy (anonymous):

ohh thx man

OpenStudy (jdoe0001):

\(\begin{array}{ccccll} 1&2&3&4\\ \hline\\ 24&24.38&24.76&25.14 \end{array}\) notice the difference between terms

OpenStudy (anonymous):

so 8 will be the next?

OpenStudy (anonymous):

yes i do

OpenStudy (jdoe0001):

in the 2, 4, 6 , yes, 8 would be next

OpenStudy (anonymous):

i meant 8 from the other

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so its A

OpenStudy (jdoe0001):

yes \(\begin{array}{ccccll} 1&2&3&4\\ \hline\\ 24&24.38&24.76&25.14\\ \hline\\ &24+0.38&24.38+0.38&24.76+0.38 \end{array} \)

OpenStudy (anonymous):

got it thx

OpenStudy (jdoe0001):

yw